OpenStudy (jenny.andrade):

Determine the number of solutions the system of 1/3 y - 5 = 2x and y = 3(2x + 5) will have.

OpenStudy (jmark):

Given 1/3 y - 5 = 2x y-15=6x........(1) and y = 3(2x + 5) y=6x+15......(2) substitute y value from equation (2) in equation (1) 6x+15-15=6x 0=0 Hence system have no solution Syllabus of ICSE class 8 - http://icse.edurite.com/icse-syllabus/icse-syllabus-for-class-8.html